OpenStudy (anonymous):

solve for right triangle ABC with the given parts a=6, b=7

OpenStudy (fibonaccichick666):

Can you draw a picture to start please?

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

Good, so now, what theorem do we know about right triangles?

OpenStudy (fibonaccichick666):

Hint: it is very famous

OpenStudy (anonymous):

the Pythagorean theorem

OpenStudy (fibonaccichick666):

good and what is the pythagorean theorem?

OpenStudy (anonymous):

\[a ^{2} + b ^{2} =c ^{2}\]

OpenStudy (fibonaccichick666):

alright, and what does c represent?

OpenStudy (anonymous):

the hypotenuse

OpenStudy (fibonaccichick666):

good so now, can you input your values into the pythagorean thm?

OpenStudy (anonymous):

\[7^{2} + 6^{2} = c ^{2}\]

OpenStudy (fibonaccichick666):

good, so now, can you simplify the left side?

OpenStudy (anonymous):

49+36=[c ^{2}\]

OpenStudy (fibonaccichick666):

good good now combine like terms

OpenStudy (anonymous):

\[85=c ^{2}\]

OpenStudy (anonymous):

\[\sqrt{85}=\sqrt{c}\]

OpenStudy (fibonaccichick666):

check that last step

OpenStudy (anonymous):

\[\sqrt{85}=C ^{2} \] ?

OpenStudy (anonymous):

that one?

OpenStudy (fibonaccichick666):

Yea, that step is not correct

OpenStudy (fibonaccichick666):

so you are taking the square root of both sides right?

OpenStudy (anonymous):

yeah

OpenStudy (fibonaccichick666):

so can you tell me what the square root of something squared is?

OpenStudy (anonymous):

i dont know but i thought it would cancel out as in \[\sqrt{85} = \sqrt{C ^{2}}\]

OpenStudy (anonymous):

and then \[\sqrt{85} = C\]

OpenStudy (fibonaccichick666):

yep, that is right now

OpenStudy (anonymous):

but how can i get \[\sqrt{85}\]

OpenStudy (fibonaccichick666):

oh that's the answer, unless you want to try and simplify it, but I'd just leave that as is

OpenStudy (anonymous):

oh ok so then am i supposed to solve for the angles

OpenStudy (fibonaccichick666):

Well, I mean you could if you have learned the trig functions, but usually it just means to find all of the sides

OpenStudy (anonymous):

yeah i did but im confused on how to solve for the angles

OpenStudy (fibonaccichick666):

well, what is the definition of tangent?

OpenStudy (anonymous):

opposite over adjacent

OpenStudy (fibonaccichick666):

good so label the angles A and B, what is tan A? tanB?

OpenStudy (anonymous):

but how will i know what is the opposite or adjacent when im not given any other number

OpenStudy (fibonaccichick666):

oh just label whichever you want A and the other B, that's all the info you require

OpenStudy (anonymous):

oh ok

OpenStudy (fibonaccichick666):

did you get it?

OpenStudy (anonymous):

lol....no

OpenStudy (fibonaccichick666):

show me your work?

OpenStudy (anonymous):